# On Singular Integrals and Quantitative Rectifiability in Parabolic Space and the Heisenberg Group

### v1

@@ 7. Examples
    Suppose $\nu$ is a measure on $(\mathbb{R},\sqrt{\,\cdot\,})$ satisfying
@@
    Suppose that $L$ is a $(k+2)$-plane in $\mathbb{R}^{n}$, and and if $\mu=\mathcal{H}^{k+2}_{|L}\times\delta_{0}$ (where $\delta_{0}$ is the Dirac measure on $\mathbb{R}$), then $\mu$ is a $(k+2)$-regular measure for which all CZOs are bounded in $L^{2}(\mu)$.
The measure $\mu$ in parts (A), (B), and (C) are parabolic vampiric measures. In fact, the following shows that any parabolic $(k+2)$-ADR vampiric measures with a product structure is bounded in $L^{2}(\mu)$. This result can be shown without the product structure, but our proof is long, and so we omit it here. Notice that part (C) is just a classical result about singular integrals in $\mathbb{R}^{k+2}$, so we shall concentrate on (A) and (B).
To prove this result we shalll verify the T(1)-theorem in spaces of homogeneous type by David-Journé-Semmes, as presented in the book Deng-Han [DH].
@@
$\langle T(\varphi,\psi\rangle_{\mu}=\frac{1}{2}\iint K(X-Y,t-s)H(X,Y,t,s)d\sigma(Y)d\sigma(X)d\nu(s)d\nu(t)$
@@
$H(X,Y,t,s)=\frac{1}{2}\Bigl[\varphi(Y,s)\psi(X,t)-\varphi(X,s)\psi(Y,t)\Bigl].$
@@
and therefore the $(n+1)$-regularity of $\mu$ ensures that the integral defining $\langle T\varphi,\psi\rangle_{\mu}$ is absolutely convergent.
If $R>0$ and $\varphi,\psi\in\operatorname{Lip}_{0}(B(x_{0},R))$ with $\|\varphi\|_{\operatorname{Lip}}\leq 1/R$ and $\|\psi\|_{\operatorname{Lip}}\leq 1/R$, then we moreover have the estimate $H(X,Y,t,s)\lesssim\frac{1}{R}\|(X,t)-(Y,s)\|$, and since $\mu$ is $(n+1)$-regular:
$|\langle T(\varphi),\psi\rangle|\lesssim\frac{1}{R}\iint_{B(x_{0},R)\times B(x_{0},R)}\frac{1}{\|x-y\|^{n}}d\mu(x)d\mu(y)\lesssim\mu(B(x_{0},R)).$
Consequently, the weak boundedness property holds (cf. Definition 1.15 of [DH]).
@@
The hypothesis of the homogeneous $T(1)$-theorem (Theorem 1.18 of [DH]) have been verified.

### v2

@@ 7. Examples
    Suppose $\nu$ is a measure on $(\mathbb{R},\sqrt{|\,\cdot\,|})$ satisfying
@@
    Suppose that $L$ is a $(k+2)$-plane in $\mathbb{R}^{n}$, and $\mu=\mathcal{H}^{k+2}_{|L}\times\delta_{0}$ (where $\delta_{0}$ is the Dirac measure on $\mathbb{R}$). Then $\mu$ is a $(k+2)$-regular measure for which all CZOs are bounded in $L^{2}(\mu)$.
The measure $\mu$ in parts (A), (B), and (C) are parabolic vampiric measures. In fact, the following shows that any parabolic $(k+2)$-ADR vampiric measure $\mu$ with a product structure has the property that all CZOs are bounded in $L^{2}(\mu)$. This result can be shown without the product structure, but our proof is long, and so we omit it here. Notice that part (C) is just a classical result about singular integrals in $\mathbb{R}^{k+2}$, so we shall concentrate on (A) and (B).
To prove this result we shall verify the T(1)-theorem in spaces of homogeneous type by David-Journé-Semmes, as presented in the book Deng-Han [23].
@@
$\langle T\varphi,\psi\rangle_{\mu}=\frac{1}{2}\iint K(X-Y,t-s)H(X,Y,t,s)d\sigma(Y)d\sigma(X)d\nu(s)d\nu(t)$
@@
$H(X,Y,t,s)=\frac{1}{2}\Bigl[\varphi(Y,s)\psi(X,t)-\varphi(X,t)\psi(Y,s)\Bigl].$
@@
and therefore the $(k+2)$-regularity of $\mu$ ensures that the integral defining $\langle T\varphi,\psi\rangle_{\mu}$ is absolutely convergent.
If $R>0$ and $\varphi,\psi\in\operatorname{Lip}_{0}(B(x_{0},R))$ with $\|\varphi\|_{\operatorname{Lip}}\leq 1/R$ and $\|\psi\|_{\operatorname{Lip}}\leq 1/R$, then we moreover have the estimate $H(X,Y,t,s)\lesssim\frac{1}{R}\|(X,t)-(Y,s)\|$, and since $\mu$ is $(k+2)$-regular:
$|\langle T(\varphi),\psi\rangle|\lesssim\frac{1}{R}\iint_{B(x_{0},R)\times B(x_{0},R)}\frac{1}{\|(X,t)-(Y,s)\|^{n}}d\mu(X,t)d\mu(Y,s)\lesssim\mu(B(x_{0},R)).$
Consequently, the weak boundedness property holds (cf. Definition 1.15 of [23]).
@@
The hypothesis of the homogeneous $T(1)$-theorem (Theorem 1.18 of [23]) have been verified.
Competing Interests: The authors declare that they have no financial interests that are directly or indirectly related to the work submitted for publication.